# In-context Learning in Presence of Spurious Correlations

### v1

@@ 2.2 The proposed approach of constructing ICL instances
Figure 2 compares the proposed and naive approaches with and without input dimension permutations. Without random permutations, the proposed approach outperforms the naive approach marginally. However, the same is not true with random permutations. We found that image embeddings of DINOv2 have a bias towards representing objects more than backgrounds, alleviating the challenge posed by the spuriously correlated background in Waterbirds. In fact, the linear probing accuracy of the spurious feature is just $\approx 82$%. For this reason, we create a modified version of Waterbirds by adding a constant vector $\tilde{s}$ or $-\tilde{s}$ to image embeddings based on the spurious feature $s$. We scale $\tilde{s}$ to have its norm equal to the average norm of image embeddings and verify that the linear probing accuracy of the spurious feature becomes 100%. On this modified Waterbirds dataset, which we name Waterbirds-severe, we see a large separation between the naive and proposed approaches (see Figure 3). We also see that without permutations, both naive and proposed approaches perform identically, indicating no robustness to the spurious correlation. This is expected, because in the absence of in-context learning, we can think of the naive and proposed approaches as standard and reweighted empirical risk minimization with a complex classification head, respectively. It has been observed that sample reweighting is not effective in overparameterized settings as all training examples will be perfectly fitted (Byrd and Lipton, 2019; Menon et al., 2021).
@@ 2.3 Comparison with conventional learning algorithms
Now that we have established the efficacy of the proposed technique, we compare it to a few established algorithms, such as 1-NN, ERM, and GroupDRO, that last of two have been historically hard to outperform (Gulrajani and Lopez-Paz, 2021; Koh et al., 2021). Comparing to more existing methods designed for robustness to spurious correlations is outside of the goal of this work, namely studying limits of in-context learning. In our comparisons, we follow the evaluation recipe used for the in-context learners. Namely, we evaluate each baseline on 8192 sequences by training on the context part of the sequence and making a prediction on the single query. More information about hyperparameters and model selection is presented in Appendix A.
@@
Figures 5(a) and 5(b) compare the proposed and baseline approaches on Waterbirds and Waterbirds-severe respectively. On Waterbirds, the proposed method outperforms ERM and GroupDRO on almost all context lengths, but is better than 1-NN only for short context lengths. The good performance of 1-NN is due to the bias in DINOv2 representations. On Waterbirds-severe, the proposed method outperforms the baselines at all context lengths. From these results, we conclude that this in-context learner implements none of these algorithms.
@@
It is worth noting that baseline worst-group accuracies at $n=512$ are actually higher than what we get when training on the entire dataset. For example, on Waterbirds, 1-NN gets only 90.03 % worst-group accuracy, while ERM gets 84.23 $\pm$ 0.17 % and GroupDRO gets 92.43 $\pm$ 0.24 %. This is due to balanced sampling of classes and setting the minority ratio to 10% withing each class, which is higher than the minority ratio of $\approx 5\%$ in the original Waterbirds dataset. One can think of our resampling as a weaker form of down-sampling which has been found to be helpful in presence of spurious correlations (Nagarajan et al., 2021; Menon et al., 2021; Idrissi et al., 2022).
Figure 4: Worst-group test accuracies on CelebA for the proposed approach and conventional methods such as 1-NN, ERM, and GroupDRO. Shaded regions show standard deviation across 5 training runs. The “+G” marker indicates setting $\tilde{y}_{i}$ to represent $g_{i}$, i.e., passing context examples groups as input. Please refer to Section 3 for more information on this.
Additionally, we verify our findings on another popular dataset CelebA (Liu et al., 2015) designed for blond vs non-blond person classification, with sex being a spurious variable. Unlike, Waterbirds, the spurious feature is asymmetric in CelebA, as blond and non-blond women are equally represented, while blond men are significantly infrequent compared to non-blond men. In particular, we verify the two shortcomings of the conventional approach and demonstrate the efficacy of the proposed techniques compared to the baselines (please see Table 4 and Figure 4 of Appendix B).
@@ 2.4 Generality of the learned algorithm
Figure 5: Worst-group test accuracies on Waterbirds and Waterbirds-severe for the proposed approach and conventional methods such as 1-NN, ERM, and GroupDRO. Majority-group accuracies are reported in Figure 14 of Appendix B. Shaded regions show standard deviation across 5 training runs.
Since we train in-context learners on ICL instances of a single task, a natural question arises whether the learned algorithm can generalize to unseen tasks. Without permuting input dimensions, the model does not learn to do in-context learning. Thus, we can not hope for any generality without permuting input dimensions. We take the model obtained with the “Proposed + P” technique and probe generality of its in-context learning by evaluating on various datasets. We start by swapping the labels of two classes in Waterbirds at evaluation and observe $\approx 2$ p.p. overall accuracy drop and $\approx 5$ p.p. worst-group accuracy drop. Despite the worsened performance, this indicates that the model treats class labels symbolically, which is remarkable given that the semantics of labels were constant during training. However, when we evaluate on Waterbirds-severe, it gets 100% accuracy on the majority groups and 0% accuracy on minority groups. Additionally, when we switch the task to predicting the background in the original Waterbirds dataset (now the class becomes a spurious feature), the overall test accuracy drops to 54.4%, while the worst-group accuracy drops to 9.3%.
@@
These OOD evaluation results indicate that the learned algorithm does something specific to the spurious feature of Waterbirds. We hypothesize that it learns to ignore this particular spurious feature. To test this, we evaluate on group-balanced Waterbirds sequences, with the task set to predicting background, and get 58.5% overall accuracy and 41.3% worst-group accuracy. One potential way of improving generality and possibly also performance, is passing example groups as input, i.e., setting $\tilde{y}_{i}$ to represent $g_{i}$. We did not observe performance improvements and increase of generality of the learned algorithm when passing groups as input (see the complete results in Tables 1 and 2 of Appendix B). Thus, we conclude that when all ICL instances are derived from the task, the learned algorithm is inherently tied to the spurious feature of that task.

### v2

@@ 2.2 The proposed approach of constructing ICL instances
Figure˜2(a) compares the proposed and naive approaches with and without input dimension permutations. Without random permutations, the proposed approach marginally outperforms the naive approach. However, the same is not true with random permutations. We found that image embeddings of DINOv2 have a bias toward representing objects more than backgrounds, alleviating the challenge posed by the spuriously correlated background in Waterbirds. In fact, with 512 training examples, the linear probing accuracy of the spurious feature is only $\approx 65$%, while that of the label is $\approx 95$%. For comparison, with ResNet-50 (He et al., 2016) embeddings, the linear probing accuracies of the spurious feature and label are $\approx 81$% and $\approx 85$% respectively.
For this reason, we create a modified version of Waterbirds by adding a constant vector $\tilde{s}$ or $\tilde{s}$ to image embeddings based on the spurious feature $s$. We scale $\tilde{s}$ to have its norm equal to the average norm of image embeddings and verify that the linear probing accuracy of the spurious feature becomes 100%. On this modified Waterbirds dataset, which we call Waterbirds-severe, we see a large separation between the naive and proposed approaches (see Figure˜2(b)). We also see that without permutations, both the naive and proposed approaches perform identically, indicating no robustness to the spurious correlation. This is expected, because in the absence of in-context learning, we can think of the naive and proposed approaches as standard and reweighted empirical risk minimization with a complex classification head, respectively. Sample reweighting has been observed to be ineffective in overparameterized settings, as all training examples will be perfectly fitted (Byrd & Lipton, 2019; Menon et al., 2021).
@@ 2.3 Comparison with conventional learning algorithms
Now that we have established the efficacy of the proposed technique, we compare it to a few established algorithms, such as 1-NN, ERM, and GroupDRO, of which the latter two have been historically hard to outperform (Gulrajani & Lopez-Paz, 2021; Koh et al., 2021). A comparison to more methods designed for robustness to spurious correlations is outside of the goal of this work, namely, studying limits of in-context learning. In our comparisons, we follow the evaluation recipe used for the in-context learners. We evaluate each baseline on 8192 sequences by training on the context part of the sequence and making a prediction on the single query. More information about hyperparameters and model selection is presented in Appendix˜A.
@@
Figures˜3(a) and 3(b) compare the proposed and baseline approaches on Waterbirds and Waterbirds-severe respectively. On Waterbirds, the proposed method outperforms ERM and GroupDRO on almost all context lengths, but is better than 1-NN only for short context lengths. The good performance of 1-NN is due to the bias in DINOv2 representations. On Waterbirds-severe, the proposed method outperforms the baselines at all context lengths. From these results, we conclude that in-context learners obtained with the proposed approach implement none of these algorithms.
@@
It should be noted that worst-group accuracies of baselines at $n=512$ are actually higher than what we get when training on the entire dataset. For example, on Waterbirds, 1-NN gets only 90.03 % worst-group accuracy, while ERM gets 84.23 $\pm$ 0.17 % and GroupDRO gets 92.43 $\pm$ 0.24 %. This is due to balanced class sampling and setting the minority ratio to 10% within each class, which is higher than the minority ratio of $\approx 5\%$ in the original Waterbirds dataset. One can think of our resampling as a weaker form of down-sampling which has been found to be helpful in presence of spurious correlations (Nagarajan et al., 2021; Menon et al., 2021; Idrissi et al., 2022).
Additionally, we verify our findings on one more image classification task CelebA (Liu et al., 2015) and on a natural language inference task MultiNLI (Williams et al., 2018a). CelebA is blond vs non-blond person classification, with sex being a spurious variable. Unlike Waterbirds, the spurious feature is asymmetric in CelebA, as blond and non-blond women are equally represented, while blond men are significantly infrequent compared to non-blond men. In MultiNLI, given a pair of sentences, a premise and a hypothesis, the task is to determine whether the hypothesis is entailed by, neutral with, or contradicts the premise. Prior work has observed that there is a spurious correlation between contradictions and the presence of the negation words nobody, no, never, and nothing (Gururangan et al., 2018). For our experiments, we frame a binary classification task entails or neutral vs. contradicts, with a binary spurious feature, which is 1 if and only if the hypothesis has one of the four negation words. For both datasets, we verify the two shortcomings of the conventional approach and demonstrate the efficacy of the proposed techniques compared to the baselines. The results for CelebA are presented in Table˜4 and Figure˜20(a), while the results for MultiNLI are presented in Table˜6 and Figure˜20(b).
@@ 2.4 Generality of the learned algorithm
Figure 3: Worst-group test accuracies on Waterbirds and Waterbirds-severe for the proposed approach and conventional methods such as 1-NN, ERM, and GroupDRO. Majority-group accuracies are reported in Figure˜18 of Appendix˜C.
Since we train in-context learners on ICL instances of a single task, a natural question arises whether the learned algorithm can generalize to unseen tasks. Without permuting input dimensions, the model does not learn to do in-context learning. Thus, we cannot expect any generality without permuting input dimensions. We take the model obtained with the “Proposed + P” approach and probe its in-context learning generality by evaluating on various tasks. We start by swapping the labels of two classes in Waterbirds at evaluation and observe $\approx 2$ p.p. overall accuracy drop and $\approx 5$ p.p. worst-group accuracy drop. Despite the worsened performance, this indicates that the model treats class labels symbolically, which is remarkable as labels had consistent semantics during training. However, when we evaluate on Waterbirds-severe, it gets 100% accuracy on the majority groups and 0% accuracy on minority groups. Additionally, when we switch the task to predicting the background in the original Waterbirds dataset (now the class becomes a spurious feature), the overall test accuracy drops to 54.4%, while the worst-group accuracy drops to 9.3%.
@@
Figure 4: Linear probing accuracy of the background variable at various layers of in-context learner transformers trained on Waterbirds.
These OOD evaluation results indicate that the learned algorithm does something specific to the spurious feature of Waterbirds. We hypothesize that it learns to ignore this particular spurious feature. To test this, we evaluate on group-balanced Waterbirds sequences, with the task set to predicting background, and get 58.5% overall accuracy and 41.3% worst-group accuracy. Additionally, we do a forward pass on 1024 ICL sequences and collect final query representations at various layers of the transformer. We then do a linear probing (512 examples for probe training and 512 for validation) to measure predictability of the background variable. We find that the “Proposed + P” approach reduces background information effectively as we sweep from input to the final layer, while the “Naive” fails to reduce background probing accuracy (see Figure˜4).
One potential way of improving generality and possibly also performance is passing example groups as input, i.e., setting $\tilde{y}_{i}$ to represent $g_{i}$. We did not observe performance improvements or an increase in generality of the learned algorithm when passing groups as input (see Tables˜1 and 2 of Appendix˜C). Thus, we conclude that when all ICL instances are derived from a single task, the learned algorithm is inherently tied to the spurious feature of that task.
@@ 3 In-context learning based on a diverse set of tasks
In Section˜2, we showed that it is possible to obtain a good in-context learner for a given task, but it fails to generalize to tasks with different spurious features. A better in-context learner should detect spurious features from context and make predictions without employing them. In this section, we explore the possibility of obtaining such a learner by training on a diverse set of ICL tasks. Since there exist few suitable datasets, we synthesize binary classification tasks with a single binary spurious feature, aiming to capture “structure” present in existing datasets. In short, given a standard binary classification task, say cat vs. dog classification, for a sampled minority of cats we overwrite some of their features with those of random dogs. Similarly, we do an analogous operation for a sampled minority of dogs. In this way some cats share dog features and vice versa. To create a diverse pool of in-context learning instances, we vary the two classes and the subset of grafted features. Please refer to Figure˜16 for an illustration of this grafting operation.
More concretely, we consider the iNaturalist dataset (Van Horn et al., 2018), which contains images from 5,089 natural fine-grained categories and filter out categories that have fewer than 500 images. For testing purposes, from remaining 239 categories we set apart the ones belonging to Amphibia and Mammalia supercategories, along with 10% of random categories. We denote the set of these 48 categories as $\mathcal{C}_{\mathrm{ood}}$, and the set of remaining 191 categories as $\mathcal{C}_{\mathrm{id}}$, which we use to create ICL instances for training. For each category in $\mathcal{C}_{\mathrm{id}}$, we hold out half of the examples as in-distribution validation set. To generate an ICL instance, we randomly sample two distinct classes from $\mathcal{C}_{\mathrm{id}}$ and sample $n/2$ images from the training split of each class uniformly at random without replacement. Please refer to Figure˜15 for an illustration of our preprocessing of iNaturalist. We then do the grafting operation, setting the minority group ratio within each class to 10%. We select the grafted features randomly, by first picking a subset size $k$ uniformly at random from 0 to 199, and then sampling a random subset of embedding dimensions of size $k$. With this we get $n$ examples that form the context part of the instance. Abandoning the naive approach and focusing on the proposed one, for each class we sample $n/2$ queries from the remaining examples uniformly at random with replacement and do the grafting operation with 50% minority group ratio.
Figure 5: Minority-group accuracy on the OOD test set of iNaturalist for the proposed approach with or without permuting input dimensions and promoting induction heads.
Figure 6: Minority-group accuracy on the OOD test set of iNaturalist for the best proposed approach with or without passing group information as input.
Following the experiments in Section˜2, we train the same transformer with the proposed approach on 4M ICL instances with $n=400$ context examples. We use the same optimizer and sweep the learning rate in the same range, selecting the best value based on the average minority-group accuracy (defined exactly in Appendix˜C) on instances where both categories belong to $\mathcal{C}_{\mathrm{ood}}$ and thus were not observed during training. The results presented in Figure˜6 indicate a major difference compared to the results in the single-task regime, namely, the proposed approach learns to do in-context learning to some extent without permuting embedding dimensions. As expected, we see much better performance with permuted embedding dimensions. Notably, comparing majority-group and minority-group accuracies of the proposed approach with permutations (Figure˜23), we see almost no sign of reliance on spurious features.
Promoting emergence of induction heads.   In-context learning ability has been linked to induction heads, which are a specific type of circuit found within large language models that implement the operation of looking back over the sequence for finding previous instances of the current token and copying what comes after that (Olsson et al., 2022). Inspired by this, we propose a data preparation technique that promotes learning of induction heads. With probability $p$, we replace each intermediate query independently with a random example from the preceeding part of the context (see Figure˜14 for an illustration of this technique). Note that this type of “hinting” is not possible in the naive approach and is enabled by the introduction of intermediate queries. In all experiments with this technique enabled, we just set $p=0.25$. We observed that training of typical runs escapes the initial loss plateau faster with this technique (in about 3k iterations compared to about 10k iterations). Moreover, we see modest performance gains in iNaturalist experiments (see Figure˜6, where +I stands for this technique).
Figure 7: Minority-group accuracy on the OOD test set of iNaturalist for the best variant of proposed approach and conventional methods such as 1-NN, ERM, and GroupDRO.
Figure 8: Worst-group test accuracy on Waterbirds for the best variant of proposed approach trained on iNaturalist and for methods such as 1-NN, ERM, and GroupDRO.
Passing example groups as input.   In contrast to the findings in the single-task setting of Section˜2, we observed that setting $\tilde{y}_{i}$ to represent group improves the proposed approach, even in addition of permuting input dimensions and promoting induction heads. One case of this is presented in Figure˜6, while more cases can be found in the complete results presented in Appendix˜C. For brevity, we mark passing groups as inputs with +G in figures and tables. Please see Figure˜13 for an illustration.
Comparison with conventional learning algorithms.   Similar to the experiments in Section˜2, we compare the best variant of the proposed approach (G + P + I) to 1-NN, ERM, and GroupDRO. The results presented in Figure˜8 show that the learned algorithm is on par with or outperforms the baselines starting at context length 32. The results at context lengths below 20 are not as informative, since our implementation of the grafting operation implies that no examples are grafted when there are less than 10 examples in a class.
Figure 9: Worst-group test accuracy on Waterbirds-severe for the best variant of proposed approach trained on iNaturalist and for methods such as 1-NN, ERM, and GroupDRO.
Generality of the learned algorithm.   To test the generality of the learned algorithm, we report evaluation results on Waterbirds (Figure˜8) and Waterbirds-severe (Figure˜9). We see that the learned algorithm outperforms baselines on Waterbirds and performs comparably to the model trained on Waterbirds itself. However, the learned algorithm fails completely on Waterbirds-severe, while the baselines give meaningful results starting at context length 32. We hypothesize that the challenge posed by the spurious features in Waterbirds-severe is significantly more severe compared to that in iNaturalist. By varying the norm of the added background vector, we interpolate between Waterbirds and Waterbirds-severe, and we see good generalization until the norm of the added vector is $\approx$40% of the average embedding norm (see Figure˜10). It is possible that the learned algorithm specializes to spurious features planted with the grafting operation, but fails to handle a strong, dense, additive spurious feature.
4 Related Work
In-weights vs in-context learning.
We observe two modes of learner behavior in our experiments. In the first mode, the learner acts like a standard supervised classifier, ignoring context examples. This mode appears when training on ICL instance of a single task without permuting input embedding dimensions. In the second mode, the learner does proper in-context learning. Our experiments indicate that both permuting embedding dimensions and increasing the number of training tasks are reliable ways of steering the model towards the in-context learning mode. The former is akin to the method of randomly projecting inputs proposed by Kirsch et al. (2022) for obtaining general-purpose in-context classifiers. Prior work has made a distinction between these two modes of learning, naming them in-weights and in-context learning. In particular, Chan et al. (2022) demonstrate that certain distributional properties of data, such as long-tail of class frequencies and bursty distribution of context example classes, can promote in-context learning when meta-training on few-shot classification instances. Singh et al. (2024) show that in-context learning behavior is not persistent and decays away with overtraining, indicating a trade-off between in-weights and in-context learning mechanisms. Moreover, they find that this in-context learning skill decay can be prevented by applying weight decay of embeddings and MLP layers, slowing down in-weights learning. Anand et al. (2024) make similar observations about these two modes of learning and propose active forgetting of token embeddings as an effective way of steering towards the in-context learning mode.
Many shot ICL.
One ancillary finding of this work is that transformers can be trained to do in-context learning of visual classification tasks when good image embeddings are provided. This is remarkable because the input dimensionality we considered is much higher than what was considered in the pioneering works of Garg et al. (2022) and Akyürek et al. (2023) (784 vs 20). Furthermore, we observe predictable performance gains from longer context sizes. The number of “shots” we consider (up to 512 examples) is well beyond what is typically considered in ICL works (up to a few dozen examples). Our findings are complementary to those of Agarwal et al. (2024)Jiang et al. (2024), and Li et al. (2024) who find that multimodal large language models, such as Gemini-1.5 Pro and GPT-4o, can benefit from large number of in-context demonstrations (up to 1000 demonstrations).
In-context learning and out-of-distribution generalization.
Closest to our work are the works that propose to make use of in-context learning for out-of-distribution generalization. Han et al. (2023) test multimodal large language models (MLLMs) on a variety of visual classification tasks. They propose to leverage in-context learning abilities of MLLMs to improve performance on specialized domains and on tasks with significant corruptions. However, they only consider the case where both context examples and query are from the target domain. Zhang et al. (2024) make similar observations, but additionally study robustness of in-context learning to distribution shifts, such as domain shifts, label shifts, and spurious correlations. They find that in-context learning is highly susceptible to label shifts and presence of spurious correlations. This is in agreement with the findings of (Ahuja & Lopez-Paz, 2023) and our work that in-context learners are brittle and don’t generalize under severe distribution shifts. Finally, Gupta et al. (2024) propose to address the problem of domain generalization (Muandet et al., 2013) by training an in-context learner that can take examples from a domain/environment and adapt to that domain in-context.
Task Diversity.
The important of task diversity for out-of-distribution generalization is well-studied outside of the in-context learning literature. In a transfer learning setting, Tripuraneni et al. (2020) derive theoretical results highlighting the importance of task diversity in addition to samples per task in obtaining transferable representations. Hendrycks et al. (2020) find that pretrained then fine-tuned models generalize to out-of-distribution examples better than only fine-tuned models, connecting the success to pretraining task diversity. Ramanujan et al. (2023) provide more direct evidence of this connection for image classification. Task diversity has also been one of the primary drivers of success of instruction tuning (Wei et al., 2022). Similarly, the excellent out-of-distribution generalization performance of the CLIP (Radford et al., 2021) has been attributed to its pretraining dataset diversity (Fang et al., 2022). Similar observations have been made for training in-context learners (Raventós et al., 2024; Goddard et al., 2025; Raparthy et al., 2024). Our findings of Section˜3 complement this literature and suggest that for training a general in-context learner robust to spurious features, one likely needs a dataset of diverse tasks with diverse spurious features.
5 Discussion and conclusion
We showed that it is possible to train an effective in-context learner tailored to a particular classification task with spurious features. We did this by introducing two key techniques: (a) permuting input embedding dimensions and (b) forming ICL sequences with intermediate queries simulating distribution shift. We provided evidence that the learned algorithm is highly competitive on the task it was trained on. However, we found that while it generalizes to other tasks without spurious features, it does not work for tasks with other spurious features. Understanding this failure mechanistically and exploring techniques for enabling better generalization are key future research directions.
We next explored training on synthetic ICL instances of diverse tasks and showed that it is possible to obtain an in-context learner that generalizes to unseen tasks, even with different data generating processes. We established the usefulness of two more techniques: (c) passing example groups as input and (d) promoting learning of induction heads by occasionally querying past context examples. We believe there is room for improving in-context learning via improved strategies of choosing intermediate queries and possibly optimizing worst-group loss. Understanding why the learned algorithm fails under extreme distribution shifts and why variants with permutations fail more (see Figure˜10) is an interesting question to explore. Another interesting direction to explore is to find out what exact algorithm is learned in the process of training on diverse tasks. Based on the results presented in this work, we conclude that the learned algorithm is neither 1-NN, ERM, or GroupDRO.
Our work has several limitations. First, training a transformer-based in-context learner with high-dimensional image embeddings is computationally costly (see Appendix˜A for information on compute resources), although it is faster than the baselines during inference. For this reason, we did not explore more datasets and pretrained image embeddings. We believe the main conclusions of our work will remain unchanged and provide an experiment on CelebA with a larger network in Appendix˜C. Second, we experimented with only one model size, width, and depth. Larger models might behave differently (Wei et al., 2023). Third, in our iNaturalist experiments, we considered only one “type” of spurious features. This choice is likely to have a significant effect on the learned algorithm and its generality. Future research should explore more ways of synthesizing spurious features and consider varying the severity of the challenge posed by spurious features. The latter can be done by considering multiple spurious features, introducing label imbalance, varying the magnitude of spurious correlations, and varying the margin spurious features provide.
Finally, we acknowledge that the proposed approach of training robust in-context learners requires spurious feature annotations, which are typically costly to obtain. As we have shown, this limitation can be addressed by creating synthetic data, in which case spurious annotations are readily available. At inference time, the learned algorithm does not require spurious annotations if it is trained with $\tilde{y}_{i}$ set to represent $y_{i}$ (i.e., ERM-like algorithm), but requires them when it is trained with passing example groups as input (i.e., $\tilde{y}$ set to represent $g_{i}$; GroupDRO-like algorithm). It is important to note that as we consider classification problems where the learner receives training data only from a single environment, spurious annotations are generally necessary to disambiguate core and spurious features.
Broader Impact Statement
Spurious correlations can appear in various domains, often in complex and unexpected ways. Given the increasing reliance on few-shot prompting in large language models, understanding and mitigating the negative effects of spurious correlations is important for ensuring the reliability and safety of these models. We hope that our work can lead to new methods and training strategies for making LLMs more robust to spurious correlations present in the context.
Acknowledgments
We would like to thank anonymous reviewers for valuable feedback. Samvel Karapetyan’s work is supported by Yandex Armenia Scholarship. The research was partially supported by the Higher Education and Science Committee of MESCS RA (Research project No. 25DD-1B082).
Appendix A Further experimental details
Baselines.
For empirical risk minimization as a baseline, we tune 2 hyperparameters: learning rate (0.01 or 0.001) and number of epochs (100 or 200). For GroupDRO we additionally tune its parameter that controls adaptiveness of group weights (0.01, 0.1, or 1) and we also try an optional strong L2 regularization (1.0 weight decay), as it has been observed to be useful for small datasets (Sagawa* et al., 2020).
Transformer-based methods.
In all transformer-based approaches, we train a causal decoder-only GPT-J transformer with 80M parameters that has 6 transformer layers with 8 multi-head attention, 768 model dimensionality, and 3072 hidden dimensionality. When training on iNaturalist, we add a layer normalization (Ba et al., 2016) on transformer input, as we expect input norms to change when we evaluate on Waterbirds-based datasets. The transformer input sequence in the proposed approach consists of 3 types of tokens: context image embeddings, query image embeddings, and label/group annotations. While the network can rely on positions and content to distinguish image embeddings from annotations, we found it to be helpful to encode token types explicitly. We do this by setting the first 3 dimensions of a token to be a one-hot vector representing token type (context image embedding, query image embedding, or annotation). When permuting dimensions, we do the permutation before encoding token types to keep the location of token-type information consistent. In our preliminary experiments and development, we used $n=128$ context length. Apart from improved performance, we did not observe significant qualitative differences when we switched to larger context lengths for final experiments.
Evaluation and model selection.
For all transformer-based approaches and baselines, we do a grid search to find the best combination of hyperparameters. In particular, we train each configuration with 5 different random seeds and select the one with the highest average test performance. Importantly, for baseline methods model selection is done for each context length independently, while for transformer-based methods model selection is done once with respect to the test performance at maximum context length observed during training. All evaluations are done on 8192 sequences, where the first $n$ examples are sampled from the corresponding train set while the query is sampled from the test set with a balanced group distribution. Finally, even when training transformers on permuted image embeddings, we do not apply permutations during evaluation. In all figures throughout this work, shaded regions show standard deviation across the 5 training runs.
Note that the most principled model selection approach would be selecting models based on a metric calculated on a dataset similar to the training set (e.g., a held-out part of training set), rather than the test set. For example, in the case of experiments on Waterbirds or Waterbirds-severe, the principled approach would be to select based on performance on sequences where the context part is sampled from the training set, while the final query is sampled from a held-out validation set with balanced group distribution. We tried this way of model selection and did not observe significant changes. In the case of experiments on iNaturalist, the principled approach would be to select based on performance on sequences where the context part is sampled from the training set, while the final query is sampled from the hold-out part the training set. We observed that this in-distribution metric is always around 99.5%-100%, and can be non-informative for model selection. This is a typical scenario in OOD generalization (see for example  (Gulrajani & Lopez-Paz, 2021) or (Wenzel et al., 2022)).
Definitions of metrics.
[…]
Compute resources.
Appendix B Additional illustrations
In this section, we present illustrations of the main techniques and data preparation steps involved in this work.
Figure 11: The attention mask in the proposed approach with $n=3$ context examples. Green cells indicate allowed attention pairs, i.e., $M_{i,j}=1$.
Figure˜11 plots the attention mask $M$ of (2) in the proposed approach for $n=3$.
Figure˜12 presents an illustration of the proposed technique of permuting image embedding dimensions (denoted by +P throughout the paper). Figure˜13 presents an illustration of the proposed approach with passing example groups as input (denoted by +G throughout the paper). Figure˜14 presents an illustration of the proposed approach with promoting emergence of induction heads (denoted by +I throughout the paper). Figure˜15 presents an illustration of our preprocessing of the iNaturalist dataset. Figure˜16 presents an illustration of the grafting operation for creating spurious features.
Figure 12: An illustration of the proposed technique of permuting image embedding dimensions (denoted by +P throughout the paper). Note that in each ICL instance we sample a new permutation, but the same permutation is used to permute dimensions of all image embeddings within one ICL instance. Whenever we use the proposed approach of forming ICL sequences (see Figure˜1(b)), the dimensions of intermediate queries are also permuted.
Figure 13: An illustration of the proposed approach with passing example groups as input (denoted by +G throughout the paper).
Figure 14: An illustration of the proposed approach with promoting emergence of induction heads (denoted by +I throughout the paper). Intermediate queries that are randomly selected to be one of the previous context examples are shown in green.
Figure 15: An illustration of our preprocessing of the iNaturalist dataset.